# Self-Supervised Learning-Based Path Planning and Obstacle Avoidance Using PPO and B-Splines in Unknown Environments

### v1

@@ IV Simulation Results
It is needless to say that since grids are occupied by ones and zeros, causing symmetries to occur in many samples, choosing the appropriate seed plays an important role in the success rate of trained networks.
@@
         Generate $\mathcal{A}$Rightarrow Calculate $C(\mathcal{S},\mathcal{A})$
@@
The steps followed in simulation and experimental sections are summarized in Algorithm 1. Depending on the complexity of the environment, one can adjust the path update rate. In environments with numerous obstacles, the path should be updated more frequently. In this study, the path is updated when $10\%$ of the path is traversed. The chosen vehicle for the simulation is a simple differential drive robot, which utilizes a PID controller to follow the path. The robot and the sensor parameters are represented in Table III. The PID controller coefficients are adjusted to minimize tracking error. It is essential to note that the controller can have a significant impact on the path planner’s functionality. Therefore, using a well-designed controller is vital.
@@
Wheelbase $0.15~{}m$
Wheel Radius $0.05~{}m$
Sensor Range $3~{}m$
Field of View $100~{}deg$
Radial Intervals $0.5~{}m$
Angular Intervals $20~{}deg$
To demonstrate the capabilities of the proposed algorithm, various scenarios are tested against the path planner. The obstacles’ edges in each scenario are represented by points to simulate feedback from conventional range sensors. Notably, the environment itself is not divided into grid cells. When the vehicle seeks to update its path at any given time step, it observes the circular segment ahead of it. Each circular grid cell containing a certain number of points is then marked as an obstacle. The output of this process is depicted in Fig. 1(b).
@@ IV-A Scenario 1: Environments with multiple small obstacles
In this scenario, the environment is populated with obstacles of uniform size, randomly positioned throughout. The vehicle starts from the origin and navigates towards the target piece by piece, relying on its limited field of view. The simulation concludes when the robot’s distance to the target is less than $3~{}m$. Fig. 6 illustrates how the robot can overcome non-convex obstacles. It is important to note that the level of non-convexity significantly impacts the algorithm’s performance. Specifically, if addressing non-convex situations requires information on past observed environments and decisions, the algorithm may fail. However, the authors have only utilized environments lacking such complexities.
@@
IV-B Scenario 2: Passing wall-like obstacles
Passing a wall-like obstacle without knowledge of its endpoint poses a significant challenge. When the vehicle encounters the wall, it must decide whether to turn right or left. Given the limited field of view, making an optimal decision regarding the distance traveled by the vehicle becomes impossible. Here, the assumption is made that the wall is short, meaning that turning right or left would not significantly impact the solution. Fig. 7 illustrates the performance of the algorithm in such scenarios.
@@ V Experimental Results
To test the algorithm performance in real situations, the Jetracer platform has been employed. The JetRacer robot, shown in Fig. 8, is a small-scale, autonomous vehicle designed for experimentation and research purposes. Equipped with a variety of sensors and a Jetson Nano onboard computer, the JetRacer is capable of executing complex navigation tasks in dynamic environments.
@@
To estimate its location, the JetRacer utilizes a combination of an Inertial Measurement Unit (IMU) and wheel odometry, which provides sufficient precision for the purposes outlined in this paper. The estimated location is then passed through an Extended Kalman Filter to reduce noise and increase the accuracy of odometry. Additionally, the JetRacer platform is equipped with a $360^{\circ}$ LIDAR sensor to detect obstacles. The LIDAR range is $12m$, but to mimic the algorithm used in this paper, the output of the LIDAR is filtered to limit the range and the field of view. It is important to note that the selection of these parameters, such as the range and field of view of the LIDAR sensor, depends on the maneuverability of the robot and the number and shape of obstacles present in the environment.
To maintain the desired path, a straightforward PID controller is utilized. The controller gains are fine-tuned through trial and error to minimize tracking errors. The vehicle maintains a constant forward speed, while the controller dynamically adjusts the steering angle based on the next waypoint provided by the path planner. Fig. 9 illustrates the interconnectedness of various components of the platform through the Robot Operating System (ROS). The entire system is designed with modularity in mind, facilitating easy replacement of the controller and path planner for future research endeavors.
@@
[…]
Figure 10: Test 1: Reaching Targets Located Behind a Wall-Shaped Obstacle
@@
In the second test, the robot is subjected to a more complex situation. Fig. 12 and Fig. 13 illustrate the results of the second test. During the test, it was observed that the field of view plays a crucial role in the vehicle’s success rate. To overcome this test, the robot required a field of view of $180^{\circ}$. It should be noted that changing the field of view is achieved simply by adjusting the filter on the LIDAR sensor, without the need to retrain the networks. In other words, the proposed method is generalizable to different environments by fine-tuning the sensor range and field of view.
Figure 12: Test 2: Robot Path and Observed Environment
@@ VI Conclusion
In this paper, we introduced SmartBSP, an AI-based path planning approach designed for autonomous robotics in complex and unknown environments. By leveraging Proximal Policy Optimization (PPO) combined with Convolutional Neural Networks (CNNs) and an Actor-Critic architecture, our method effectively processes limited LIDAR inputs to generate optimal paths in real-time. The integration of B-spline curves and a nuanced cost function accounting for path curvature, endpoint proximity, and obstacle avoidance further enhances the robustness and adaptability of the system.
Through extensive simulations and real-world experiments, we demonstrated that the proposed method can handle various challenging scenarios, including environments with multiple small obstacles and wall-like obstacles. Our results show that adjusting the field of view and sensor range significantly impacts the success rate of the vehicle, highlighting the importance of these parameters in path planning tasks. Moreover, the flexibility of the SmartBSP approach allows for easy adaptation to different environments without the need for retraining the networks, showcasing its potential for practical applications in diverse operational scenarios.
Future work will focus on further refining the algorithm to handle more complex non-convex obstacles and exploring the integration of additional sensor modalities to enhance the perception capabilities of the autonomous system. Additionally, we aim to apply the SmartBSP framework to other types of autonomous vehicles and investigate its performance in large-scale outdoor environments.

### v2

@@ IV Simulation Results
Since grids are occupied by ones and zeros, causing symmetries to occur in many samples, choosing the appropriate seed plays an important role in the success rate of trained networks.
@@
   Generate $\mathcal{A}$ $\Rightarrow$ Calculate $C(\mathcal{S},\mathcal{A})$
@@
The steps followed in simulation and experimental sections are summarized in Algorithm 1. Depending on the complexity of the environment, one can adjust the path update rate accordingly. In environments with numerous obstacles, the path should be updated more frequently. In this study, the path is updated when $10\%$ of the path is traversed. The chosen vehicle for the simulation is a simple differential drive robot which utilizes a PID controller to follow the path. The robot and the sensor parameters are represented in Table III. The PID controller coefficients are adjusted to minimize tracking error. Since the controller can significantly affect the path planner performance, using a well-designed controller is vital.
@@
Wheelbase 0.15m0.15~m
Wheel Radius 0.05m0.05~m
Sensor Range 3m3~m
Field of View 100deg100~deg
Radial Intervals 0.5m0.5~m
Angular Intervals 20deg20~deg
To demonstrate the capabilities of the proposed algorithm, various scenarios are tested against the path planner. The obstacles’ edges in each scenario are represented by points to simulate feedback from conventional range sensors. Notably, the environment itself is not divided into grid cells. When the vehicle seeks to update its path at any given time step, it observes the circular segment ahead of it. Each circular grid cell containing a certain prescribed number of points detected by LiDAR sensor is then marked as an obstacle. The output of this process is depicted in Fig. 1(b).
@@ IV-A Scenario 1: Environments with multiple small obstacles
In this scenario, the environment is populated with obstacles of uniform size, randomly positioned throughout the search area. The vehicle navigates from the origin towards the target, piece by piece, relying on its limited field of view. The simulation concludes when the robot’s distance to the target is less than a chosen threshold of 3m3~m. Fig. 6 illustrates how the robot can overcome non-convex obstacles. The level of non-convexity significantly impacts the algorithm’s performance. Specifically, if addressing non-convex situations requires information on past observed environments and decisions, the algorithm may fail. Such environments are not considered in this study.
@@
IV-B Scenario 2: Passing large obstacles
Passing a large obstacle without knowledge of its endpoint poses a significant challenge. When the vehicle encounters the wall, it must decide whether to turn right or left. Given the limited field of view, making an optimal decision regarding the distance traveled by the vehicle becomes impossible. Here, the assumption is made that the wall is short, meaning that turning right or left would not significantly impact the solution. Fig. 7 illustrates the performance of the algorithm in such scenarios.
@@ V Experimental Results
To test the algorithm performance in real situations, the Jetracer platform is employed. The JetRacer robot, shown in Fig. 8, is a small-scale, autonomous vehicle designed for experimentation and research purposes. Equipped with a variety of sensors and a Jetson Nano onboard computer, the JetRacer is capable of executing complex navigation tasks in dynamic environments.
@@
The JetRacer platform is equipped with a $360^{\circ}$ LiDAR sensor to detect obstacles. The LiDAR range is $12m$. To mimic the algorithm used in this paper, however, the LiDAR output is filtered to limit the range and the field of view to match those used in simulations. The selection of these parameters, such as the range and field of view of the LiDAR sensor, depends on the maneuverability of the robot and the number and shape of obstacles present in the environment.
To maintain the desired path, a straightforward PID controller is utilized. The vehicle maintains a constant forward speed, while the controller dynamically adjusts the steering angle based on the next waypoint provided by the path planner. Fig. 9 illustrates the interconnectedness of various components of the platform through the Robot Operating System (ROS). The entire system is designed with modularity in mind, facilitating easy substitutions of various controllers and path planners.
@@
To assess the algorithm, the vehicle is tested in two different scenarios. Fig. 10 and Fig. 11 present the results of the first test. The final goal is located behind a wall, requiring the robot to circumvent a non-convex wall to reach it. To perform this test, the vehicle’s field of view is set to $120^{\circ}$ and the sensor range to $0.8\,\text{m}$. For both experimental tests, the robot has no prior knowledge of the environment. The starting position, denoted by yellow, is considered the origin. The robot is programmed to stop when it is within a predefined threshold distance from the target position. The red points represent the observations of obstacles from the LiDAR sensor. The entire path is planned incrementally, utilizing the robot’s limited field of view.
Figure 10: Test 1: Reaching Target Located Behind a Wall-Shaped Obstacle
@@
In the second test, the robot is subjected to a more complex situation. Fig. 12 and Fig. 13 illustrate the results of the second test. During the test, it is observed that the field of view plays a crucial role in the vehicle’s success rate. To ensure adequate performance, the robot requires a field of view of $180^{\circ}$. Here, changing the field of view is achieved simply by adjusting the filter on the LiDAR sensor, without the need to retrain the networks. In other words, the proposed method is generalizable to different environments by appropriately selecting the sensor range and field of view.
Figure 12: Test 2: Avoiding non-convex obstacle
@@ VI Conclusion
In this paper, the authors introduce SmartBSP, an AI-based path planning approach designed for autonomous robotics in complex and unknown environments. By leveraging Proximal Policy Optimization (PPO) combined with Convolutional Neural Networks (CNNs) and an Actor-Critic architecture, the proposed method effectively processes limited LIDAR inputs to generate optimal paths in real-time. The integration of B-spline curves and a nuanced cost function accounting for path curvature, endpoint proximity, and obstacle avoidance further enhances the robustness and adaptability of the system.
Through extensive analytical simulations and lab experiments, the authors demonstrate that the proposed method maintains good performance in spite of challenging scenarios, including environments with multiple small obstacles and large obstacles. Experimental results show that adjusting the field of view and sensor range significantly impacts the success rate of the vehicle, highlighting the importance of these parameters in path planning tasks. Moreover, the flexibility of the SmartBSP approach allows for easy adaptation to different environments without the need for retraining the networks, showcasing its potential for practical applications in diverse operational scenarios.
Future work will focus on further refining the algorithm to take into account more complex non-convex obstacles and exploring the integration of additional sensor modalities to enhance the perception capabilities of the autonomous system. Additionally, the authors aim to apply the SmartBSP framework to other types of autonomous vehicles and investigate its performance in large-scale outdoor environments.